# COBRA: A Continual Learning Approach to Vision-Brain Understanding

### v1

@@ 6.1 Explanation of Spatial Issue in 1D fMRI Signals
In this section, We provide additional details about the spatial issue discussed in Section 3.2 regarding fMRI signals represented in 1D form. Specifically, for a given subject’s fMRI signals, we mask out all samples by setting their values to zero, retaining only 1000 continuous samples and assigning their values to one. These modified signals are then projected onto the fsaverage surface for visualization, as shown in Fig. 7. It is evident that in the 1D form, the 1000 continuous samples share strong correlations. However, when these signals are represented in 2D, the samples appear sparsely distributed. For that reason, analyzing fMRI signals in 1D form leads to a significant loss of spatial information.
6.2 Accuracy of SC and PSS Modules
In the proposed COBRA, the SC module is designed to remain fixed when training on new subjects, as it captures common vision-brain patterns learned from previous subjects. To verify this, we trained the SC module on subjects (3,4,6,8) and tested it on the unseen subjects (1,2,5,7). The SC module achieved an F1 score of 0.84, supporting our hypothesis. For the PSS module, we evaluated the PSS module’s accuracy in learning subject-specific features for each subject. It achieved a high accuracy of 98.5%, indicating that the module effectively selects prompts that represent subject-specific features. These results are presented in Table 7.
| Module | Metric | Accuracy |
| SC | f1 score | 0.84 |
| PSS | accuracy | 98.5 |
6.3 Visualization of Subject-Specific Features
In this section, we analyze the discrimination of subject-specific features, i.e., , that is extracted from the PSS module. We employ T-SNE [71] to reduce the dimension of the features. The visualization is illustrated in Fig. 8. The result demonstrates that the PSS module was able to learn the specific features that are discriminative for the subjects.
6.4 Qualitative Comparision With Previous Continual Learning Approaches
In this section, we compare the qualitative vision-brain reconstruction results of the model under four scenarios: without any continual learning (CL) approach, using LwF [30], using PLOP [12], and our proposed method, COBRA. The results are illustrated in Fig. 9 and Fig. 10. Our proposed method demonstrates significantly better and clearer reconstruction results compared to previous methods. In contrast, LwF [30] and PLOP [12] lose critical contextual details. Notably, without applying any CL approach, the generated images are incoherent and fail to reconstruct the context meaningfully.

### v3

@@ 7.2 Explanation of Spatial Issue in 1D fMRI Signals
In this section, We provide additional details about the spatial issue discussed in Section 3.2 regarding fMRI signals represented in 1D form. Specifically, for a given subject’s fMRI signals, we mask out all samples by setting their values to zero, retaining only 1000 continuous samples and assigning their values to one. These modified signals are then projected onto the fsaverage surface for visualization, as shown in Fig. 11. It is evident that in the 1D form, the 1000 continuous samples share strong correlations. However, when these signals are represented in 2D, the samples appear sparsely distributed. For this reason, analyzing fMRI signals in 1D form results in a significant loss of spatial information.